Grade 11Physical Chemistry

Calculate the value of propotionality constant k for 1 mole of a gas which occupies 22.4L of volume under the given conditions of temperature and pressure.

Rakesh
7 Years agoGrade 11

1 Answer

The Hounds Of Justice
ApprovedApproved Tutor Answer7 Years ago
As n=1 mol
     V=22.4 litre
and we know that v=nk (k is a constant)
then k=v/n = 22.4 L mol-1
Hence it is the propotionality constant